OpenStudy (anonymous):

Solve the system by elimination (Need to show work) {14x-35=7y {-25-6x=5y

OpenStudy (jdoe0001):

first of all you may want to organize the variables vertically x + y = x + y = then multiply one by a factor to cancel out or "eliminate" one

OpenStudy (anonymous):

Can someone show me how to do this

OpenStudy (jdoe0001):

well... firstlyi arrange your variables.. so use linear simplification to move them about notice how they're not aligned vertically, the "x"'s should all be in one-vertical-line same for the "y"'s preferably on the left-hand-side \(\bf 14{\color{brown}{ x}}-35=7{\color{blue}{ y}}\\ -25-6{\color{brown}{ x}}=5{\color{blue}{ y}}\)

OpenStudy (anonymous):

like this? 14x-35=7y -6x-25=5y

OpenStudy (jdoe0001):

well... that'd work, yes so let us use a factor to eliminate one one sec

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} 14x-35=7y&{\color{brown}{ \times -5}} &\to-70x+175=\cancel{ -35y }\\ -6x-25=5y&{\color{brown}{ \times 7}} &\to-42x-175=\cancel{ 35y } \\\hline\\ &&\square ?\qquad +\square ?\qquad =\square ? \end{array}\) notice that we used -5 and 7 and notice that the multipiied version resulted in a -35y and a +35y now add the rest vertically, see what you get

OpenStudy (anonymous):

8x-(-60)

OpenStudy (jdoe0001):

-(-60)?

OpenStudy (anonymous):

wouldn't it be negative or no

OpenStudy (texaschic101):

you should have reduced the first equation...makes it a lot easier

OpenStudy (jdoe0001):

well... youu'd be adding vertically, thus why aligning them vertically

OpenStudy (anonymous):

so was I wrong?

OpenStudy (jdoe0001):

well.... notice the addition you need to perform \(\begin{array}{llll} 14x-35=7y&{\color{brown}{ \times -5}} &\to-70x+175=\cancel{ -35y }\\ -6x-25=5y&{\color{brown}{ \times 7}} &\to-42x-175=\cancel{ 35y } \\\hline\\ &&\square ?\qquad +\square ?\qquad =\square ?\\ &&\qquad \Uparrow\\ &&sum\ up \end{array}\)

OpenStudy (jdoe0001):

do you see how we get the 2nd set of equations?

OpenStudy (jdoe0001):

well... do you see how we get the 2nd set of equations? do you even see the 2nd set of equations ?

OpenStudy (anonymous):

ok we multiply it by -5 and 7

OpenStudy (jdoe0001):

yeap..... and then we add vertically

OpenStudy (jdoe0001):

in the addition, notice -35y+ 35y = 0 thuse they get ELIMIINATEd

OpenStudy (anonymous):

wouldn't 175 cancel itself out because of the plus and minus sign?

OpenStudy (jdoe0001):

in this case, yes, it does so... what are you left with then \(\begin{array}{llll} 14x-35=7y&{\color{brown}{ \times -5}} &\to-70x+175=\cancel{ -35y }\\ -6x-25=5y&{\color{brown}{ \times 7}} &\to-42x-175=\cancel{ 35y } \\\hline\\ &&\square ?\qquad +0\qquad =0 \end{array}\)

OpenStudy (anonymous):

so -70+-42= -114x

OpenStudy (jdoe0001):

well... -70-42 = -112 but yes.... then that's the equation you're left with thus \(\bf \begin{array}{llll} 14x-35=7y&{\color{brown}{ \times -5}} &\to-70x+175=\cancel{ -35y }\\ -6x-25=5y&{\color{brown}{ \times 7}} &\to-42x-175=\cancel{ 35y } \\\hline\\ &&-112x\qquad +0\qquad =0 \end{array} \\ \quad \\ {\color{blue}{ -112x=0\implies x=\cfrac{0}{-112}\implies x=0}}\)

OpenStudy (jdoe0001):

so now you know what "x" is then you can plug it either equation, to get "y" by solving for "y" :) see by multiplying like so by some factor, so that it yields the same product atop and below one of the variables get ELIMINATED, thus the "elimination" method

OpenStudy (anonymous):

then to get y I just put the value of x in the original equation

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

weird how my book shows a more complex way

OpenStudy (anonymous):

thanks

OpenStudy (jdoe0001):

yw